OpenStudy (anonymous):

If the area of triangle DEF is 220.5 cm^2, what is the area of triangle ABC?

OpenStudy (anonymous):



OpenStudy (anonymous):

ABC~DEF

OpenStudy (anonymous):

from the triangle DEF you know its area and base then you can find the height

OpenStudy (anonymous):

h=2*220.5/21 =441/21 =21

OpenStudy (caozeyuan):

yes but no need

OpenStudy (caozeyuan):

for two similiar triangles, ratio of areas = ratio of sides squared

OpenStudy (anonymous):

25/49?

OpenStudy (anonymous):

You there @caozeyuan

OpenStudy (jonnyvonny):

@MaryLovell15 Still need help with it?

OpenStudy (anonymous):

Yes! :( @JonnyVonny

OpenStudy (jonnyvonny):

Mmk, so, you found the height of one triangle, and you know its base.

OpenStudy (anonymous):

Right.

OpenStudy (jonnyvonny):

So, the height and base are both 21?...

OpenStudy (jonnyvonny):

Yup, which, some triangles are like that. They are called equilateral triangles; the sides and angles are all the same.

OpenStudy (jonnyvonny):

So, you have the height and base, and you know they're similar, so you can set up a ratio between the two:

OpenStudy (jonnyvonny):

height 21 x height base 21 15 height

OpenStudy (jonnyvonny):

Are you following me? Or are you confused?

OpenStudy (anonymous):

I understand so far.

OpenStudy (jonnyvonny):

So, now that we have those ratios, we can cross multiply and solve for x. Tell me what you get and I'll tell you what I got.

OpenStudy (anonymous):

x=15

OpenStudy (anonymous):

@JonnyVonny

OpenStudy (jonnyvonny):

Yes, that is what i got aswell

OpenStudy (anonymous):

So the area is 112.5?

OpenStudy (jonnyvonny):

So, now, just get the area via: A=.5b*h

OpenStudy (jonnyvonny):

Yup, well, thats what we both got, at least.

OpenStudy (jonnyvonny):

There shouldn't be a reason why its wrong. Math does have a way at getting at you, however.:b

OpenStudy (anonymous):

I honestly don't care if it's wrong at this point. Lol, I've been stuck on this for awhile.

OpenStudy (jonnyvonny):

Lol, I was just playing around; this is the way you're suppose to do it; find comparable numbers, then cross multiply. Know this process; you'll be using it a lot.

OpenStudy (anonymous):

Thanks for your help!